capacity of a spherical capacitor is C1 when inner sphere is charged and outer sphere is earthed and C2 when inner sphere is earthed and outer sphere is charged.then C1/C2 is
(''a'' is radius of inner sphere
''b'' is radius of outer sphere)

Grade:12

2 Answers

Arun Kumar IIT Delhi
9 years ago
Here capacitance will be inversely proportional to radius =>
C1/C2=b/a.
Thanks & Regards
Arun Kumar
IIT Delhi
Askiitians Faculty
Anwesha Dash
7 years ago
When inner sphere is charged and outer sphere is earthed . The capacity of the capacitor can be calculated by potential of inner sphere calculation.C1= 4 pi £ ab/(b-a)When outer sphere is charged and inner sphere is earthEd. Let charge Q be given to the outer sphere . Q1 Charge is on outer side and Q2 is on inner side of the outer sphere. Such that Q1 + Q2 = Q -Q2 is induced on the outer side of inner sphere. Here the system is equivalent to a spherical capacitor of inner and outer radii and a spherical conductor of outer radii in parallel. So total capacitance is the sum of their individual capacitance.C2 =C1 + Capacitance of a spherical conductor of radii b= 4 pi £ ab/b-a + 4 pi £ bC2 = 4 pi £ b^2/(b-a)C1/C2 = a/b P.S Capacitance of a spherical conductor can be calculated by considering any charge given to a conductor spreads on the outer surface of it. And potential is calculated assuming the charge to be present in the centre of the conductor so that.V= q/4 pi £ bC=q/V = 4 pi £ b